# Global well-posedness of the elastic-viscous-plastic sea-ice model with the inviscid Voigt-regularisation

### 06 Mar 2026 v2

@@ 3.1 The Galerkin approximation scheme
$\partial_{t}u_{N}=\nabla\cdot\sigma_{N}+\mathbf{P}_{N}\mathcal{T}_{a}+\mathbf{P}_{N}\mathcal{T}_{w}^{N}+\Omega u^{\perp}_{N}-g\mathbf{P}_{N}\nabla H_{0},$ (25a)
$\begin{gathered}\frac{1}{\mathcal{E}}\partial_{t}\big(\sigma_{N}-\alpha^{2}\Delta\sigma_{N}\big)+\mathbf{Q}_{N}\bigg[\frac{4\mathcal{D}_{\epsilon}^{N}}{P}(\sigma_{N}-\frac{1}{2}\operatorname{Tr}\sigma_{N}\mathbb{I}_{2})\bigg]+\mathbf{Q}_{N}\bigg[\frac{\mathcal{D}_{\epsilon}^{N}}{2P}\operatorname{Tr}\sigma_{N}\mathbb{I}_{2}\bigg]\\ +\frac{\mathbf{Q}_{N}\mathcal{D}_{\epsilon}^{N}}{2}\mathbb{I}_{2}=D(u_{N}),\end{gathered}$ (25d)
$u\lvert_{t=0}=\mathbf{P}_{N}u_{0},\quad\sigma\lvert_{t=0}=\mathbf{Q}_{N}\sigma_{0},$ (25e)
where $\mathbf{P}_{N}$ is the projection of two-dimensional vector fields onto the Fourier modes up to order $N$, including the zeroth mode, while $\mathbf{Q}_{N}$ is the projection of $2\times 2$ (symmetric) matrices onto the Fourier modes up to order $N$. In system (25), we have used the notation
$\mathcal{T}_{w}^{N}$ $=c_{w}\rho_{w}\lvert U_{w}-u_{N}\rvert\bigg[(U_{w}-u_{N})\cos\theta+(U_{w}-u_{N})^{\perp}\sin\theta\bigg]$ (26)
@@
$\mathcal{D}_{\epsilon}^{N}$ $=\sqrt{\lvert D(u_{N})\rvert^{2}+\epsilon^{2}}.$ (27)
Notice that system (25) is a first-order system of ordinary differential equations for $\{\hat{u}_{k},\hat{\sigma}_{k}\}_{|k|\leq N}$, with a locally Lipschitz continuous vector field with respect to the unknowns. Thus, the Picard theory (see for example Ref. [28]) implies that, for each initial datum $(\mathbf{P}_{N}u_{0},\mathbf{Q}_{N}\sigma_{0})$, there exists a unique local-in-time solution.
@@ 3.2 Uniform-in-$(N,\epsilon)$ estimates
To establish the uniform-in-$N$ estimates for the solutions of system (25), let
$\tau_{N}\coloneqq\sigma_{N}+\frac{P}{2}\mathbb{I}_{2}.$ (28)
Recall that the pressure $P>0$ is constant. We then rewrite Eq. (25d) as follows:
$\begin{gathered}\frac{1}{\mathcal{E}}\partial_{t}\bigg(\tau_{N}-\alpha^{2}\Delta\tau_{N}\bigg)+\mathbf{Q}_{N}\bigg[\frac{4\mathcal{D}_{\epsilon}^{N}}{P}\bigg(\tau_{N}-\frac{1}{2}\operatorname{Tr}\tau_{N}\mathbb{I}_{2}\bigg)\bigg]\\ +\mathbf{Q}_{N}\bigg[\frac{\mathcal{D}_{\epsilon}^{N}}{2P}\operatorname{Tr}\tau_{N}\mathbb{I}_{2}\bigg]=D(u_{N}).\end{gathered}$ (29)
$L^{2}$-estimate. We start by computing the $L^{2}(\mathbb{T}^{2})$ estimates for the solution of system (25) on its maximal (positive) time interval of existence. Taking the $L^{2}$-inner product of (25a) and (29) with $u_{N}$ and $\tau_{N}$, respectively, applying integration by parts, and summing the results yields
$\begin{gathered}\frac{1}{2}\frac{\rm d}{\ {\rm d}t}\bigg[\lVert u_{N}\rVert_{L^{2}}^{2}+\mathcal{E}^{-1}\|\tau_{N}\|_{L^{2}}^{2}+\alpha^{2}\mathcal{E}^{-1}\|\nabla\tau_{N}\|_{L^{2}}^{2}\biggr]\\ +\int_{\mathbb{T}^{2}}\biggl[\frac{4\mathcal{D}_{\epsilon}^{N}}{P}|\tau_{N}-\frac{1}{2}\operatorname{Tr}\tau_{N}\mathbb{I}_{2}|^{2}+\frac{\mathcal{D}_{\epsilon}^{N}}{2P}|\operatorname{Tr}\tau_{N}|^{2}\biggr]\,\ {\rm d}{x}\\ =\int_{\mathbb{T}^{2}}\bigg[\mathcal{T}_{a}+\mathcal{T}_{w}^{N}+\Omega u^{\perp}_{N}-g\nabla H_{0}\bigg]\cdot u_{N}\ {\rm d}{x}\leq C+C\|u_{N}\|_{L^{2}}^{2},\end{gathered}$ (30)
@@
where we use the fact that $\cos\theta\geq 0$ since $\theta\in[0,\frac{\pi}{4}]$. Notice that the constant $C\in(0,\infty)$ in (30) depends only on $\Omega,g,P,H_{0},U_{a},U_{w},c_{a},c_{w},\phi,\theta,\rho_{a},$ and $\rho_{w}$, but not on $\epsilon$ and $N$. We remark that estimate (30) would also hold if one considers the intermediate model with the regularised and nonsimplified strain rate $\overline{\mathcal{D}}_{\epsilon}^{N}$ instead of $\mathcal{D}_{\epsilon}^{N}$ (as both have definite sign). Moreover, we also have used the following cancellation, which is similar in structure to the Oldroyd-B model,
$\int_{\mathbb{T}^{2}}\bigg[u_{N}\cdot(\nabla\cdot\sigma_{N})+\tau_{N}:D(u_{N})\bigg]\ {\rm d}{x}=0.$ (31)
Therefore, applying the Grönwall inequality to estimate (30) on a time interval of existence $[0,T]$ of the approximating Galerkin solution leads to
$\sup_{t\in[0,T]}\bigg[\lVert u_{N}(t)\rVert_{L^{2}}^{2}+\mathcal{E}^{-1}\left\lVert\sigma_{N}(t)+\frac{P}{2}\mathbb{I}_{2}\right\rVert_{L^{2}}^{2}+\alpha^{2}\mathcal{E}^{-1}\left\lVert\nabla\bigg[\sigma_{N}(t)+\frac{P}{2}\mathbb{I}_{2}\bigg]\right\rVert_{L^{2}}^{2}\bigg]\leq Ce^{CT},$ (32)
uniformly in $N$ (and also $\epsilon$). Since the right-hand side remains finite for any $T\in(0,\infty)$, we conclude that the time interval of existence of the approximating Galerkin solutions is $0,\infty)$ and estimate (32) therefore holds for any $T\in(0,\infty)$.
@@
$V_{w}:=U_{w}-u_{N}.$ (33)
Recalling (26), one can write
$\begin{gathered}-\int_{\mathbb{T}^{2}}\mathbf{P}_{N}\mathcal{T}_{w}^{N}\cdot\Delta u_{N}\ {\rm d}{x}=\sum_{i=1}^{2}\int_{\mathbb{T}^{2}}\mathbf{P}_{N}\partial_{i}\mathcal{T}_{w}^{N}\cdot\partial_{i}u_{N}\ {\rm d}{x}=\sum_{i=1}^{2}\int_{\mathbb{T}^{2}}\partial_{i}\mathcal{T}_{w}^{N}\cdot\partial_{i}u_{N}\ {\rm d}{x}\\ =\sum_{i=1}^{2}\int_{\mathbb{T}^{2}}\partial_{i}\mathcal{T}_{w}^{N}\cdot\partial_{i}{U_{w}}\ {\rm d}{x}-\sum_{i=1}^{2}\int_{\mathbb{T}^{2}}c_{w}\rho_{w}\bigl[\partial_{i}(\lvert V_{w}\rvert V_{w})\cos\theta\bigr]\cdot\partial_{i}V_{w}\ {\rm d}{x}\\ -\sum_{i=1}^{2}\int_{\mathbb{T}^{2}}c_{w}\rho_{w}\partial_{i}(|V_{w}|V_{w}^{\perp})\cdot\partial_{i}V_{w}\sin\theta\ {\rm d}{x}=:I_{1}+I_{2}+I_{3},\end{gathered}$ (34)
@@
$\partial_{i}(|V_{w}|V_{w})=|V_{w}|\partial_{i}V_{w}+\dfrac{V_{w}\cdot\partial_{i}V_{w}}{|V_{w}|}V_{w}.$ (35)
@@
$|I_{1}|\lesssim\|V_{w}\|_{L^{2}}\|\nabla V_{w}\|_{L^{2}}\lesssim 1+\|{u_{N}}\|_{L^{2}}^{2}+\|{\nabla u_{N}}\|_{L^{2}}^{2},$ (36)
@@
$=\left\lvert\sum_{i=1}^{2}\int_{\mathbb{T}^{2}}c_{w}\rho_{w}\dfrac{V_{w}\cdot\partial_{i}V_{w}}{|V_{w}|}V_{w}^{\perp}\cdot\partial_{i}V_{w}\sin\theta\ {\rm d}{x}\right\rvert$ (37)
$\leq c_{w}\rho_{w}|\sin\theta|\int_{\mathbb{T}^{2}}|V_{w}||\nabla V_{w}|^{2}\ {\rm d}{x}.$ (38)
Meanwhile, substituting (35) in $I_{2}$, one can compute
$I_{2}=-c_{w}\rho_{w}\cos\theta\int_{\mathbb{T}^{2}}\left(|V_{w}||\nabla V_{w}|^{2}+\frac{1}{4}\frac{\left|\nabla\big(|V_{w}|^{2}\big)\right|^{2}}{|V_{w}|}\right)\ {\rm d}{x}.$ (39)
Consequently, (34)–(39) leads to
$\int_{\mathbb{T}^{2}}\mathbf{P}_{N}\mathcal{T}_{w}^{N}\cdot\Delta u_{N}\ {\rm d}{x}=\int_{\mathbb{T}^{2}}\nabla\mathcal{T}_{w}^{N}:\nabla u_{N}\lesssim 1+\|{u_{N}}\|_{L^{2}}^{2}+\|{\nabla u_{N}}\|_{L^{2}}^{2}$ (40)
@@
Taking the $L^{2}$-inner product of (25a) and (25d) with $\Delta u_{N}$ and $\Delta\sigma_{N}$, respectively, applying integration by parts, and summing up the results leads to the following estimate
@@
$+\lVert\Delta\sigma_{N}\rVert_{L^{2}}^{2}+\lVert\nabla u_{N}\rVert_{L^{2}}^{2}+\|u_{N}\|_{L^{2}}^{2}+1$ (42)
where we have used (40), the Brézis-Gallouët-Wainger inequality (16), as well as Young’s inequality. Note that we have introduced a constant $R_{1}>0$ in order to make the quantity $R_{1}\sigma_{N}$ dimensionless (so that we can apply the Brézis-Gallouët-Wainger inequality), and in particular $R_{1}\sigma_{N}$ does not have physical units. One can choose $R_{1}=\lVert\sigma_{0}\rVert_{L^{\infty}}^{-1}$ for example (for nonzero initial data). We observe that for the ${H}^{1}$-estimate one can also treat the nonsimplified strain rate $\overline{\mathcal{D}}_{\epsilon}^{N}$ instead of $\mathcal{D}_{\epsilon}^{N}$ by using estimate (79). In addition, we have used the following cancellation identity
$\sum_{i=1}^{2}\int_{\mathbb{T}^{2}}\bigg[\partial_{i}u_{N}\cdot(\nabla\cdot\partial_{i}\sigma_{N})+\partial_{i}\sigma_{N}:D(\partial_{i}u_{N})\bigg]\ {\rm d}x=0.$ (43)
Thanks to the $L^{2}$-estimate in (32), we can apply the logarithmic Grönwall inequality from Lemma 2.2 to (42), where we set
@@
$\sup_{t\in[0,T]}\biggl[\|\nabla u_{N}\|_{L^{2}}^{2}+\mathcal{E}^{-1}\|\nabla\sigma_{N}\|_{L^{2}}^{2}+\alpha^{2}\mathcal{E}^{-1}\|\Delta\sigma_{N}\|_{L^{2}}^{2}\biggr]\leq C\exp(\exp(CT)),$ (44)
for some $C$ independent of $N,\epsilon$, and $T$, but depending on $\mathcal{E},\alpha$, and the initial data. Note that the double-exponential nature of the bound comes from the application of the logarithmic Grönwall inequality to the single-exponential $L^{2}$-estimate (32).
@@
$\nabla\mathcal{D}_{\epsilon}^{N}$ $=\frac{D(u_{N}):\big[\nabla D(u_{N})\big]}{\mathcal{D}_{\epsilon}^{N}}=\frac{\frac{1}{2}\nabla\lvert D(u_{N})\rvert^{2}}{\mathcal{D}_{\epsilon}^{N}},$ (45)
@@
In addition, we recall the following computation from p. 22-23 in Ref. [57], for $i,j=1,2$, with $V_{w}$ given in (33),
$\partial^{2}_{ij}\big(\lvert V_{w}\rvert V_{w}\big)=$ $\;\lvert V_{w}\rvert\partial^{2}_{ij}V_{w}+\frac{V_{w}\cdot\partial_{i}V_{w}}{\lvert V_{w}\rvert}\partial_{j}V_{w}+\frac{V_{w}\cdot\partial_{j}V_{w}}{\lvert V_{w}\rvert}\partial_{i}V_{w}$ (46)
@@
$\lesssim\big(\lVert V_{w}\rVert_{L^{\infty}}+\lVert V_{w}\rVert_{H^{1}}\big)\lVert V_{w}\rVert_{H^{2}},$ (47)
where we have applied the Ladyzhenskaya inequality (17).
Then, after taking the $L^{2}$-inner product of (25a) and (25d) with $\Delta^{2}u_{N}$ and $\Delta^{2}\sigma_{N}$, respectively, using integration by parts, and summing up the results, one obtains
@@
$+\|\sigma_{N}\|_{H^{2}}+1\bigg]\cdot\big[\lVert u_{N}\rVert_{H^{2}}^{2}+\lVert\sigma_{N}\rVert_{H^{3}}^{2}\big]+1,$ (48)
where we have used (47), the Brézis-Gallouët-Wainger inequality (16), the Sobolev embedding theorem as well as Young’s inequality. The terms on the second line are negative definite, and in the seventh line we estimate the term $\lVert\nabla^{2}u_{N}\rVert_{L^{2}}\lVert\nabla\sigma_{N}\rVert_{L^{4}}\lVert\Delta\sigma_{N}\rVert_{L^{4}}$ by means of the Sobolev embedding theorem, which gives $\lVert u_{N}\rVert_{H^{2}}\lVert\sigma_{N}\rVert_{H^{2}}\lVert\sigma_{N}\rVert_{H^{3}}$. Applying Young’s inequality leads to some of the terms on the last line of estimate (3.2). Due to the monotonicity of the logarithm, we can replace the arguments $\lVert\sigma_{N}\rVert_{H^{2}}$ and $\lVert u_{N}\rVert_{H^{2}}$ by their sum, which facilitates the application of the logarithmic Grönwall inequality.
@@
$\int_{\mathbb{T}^{2}}\bigg[\Delta u_{N}\cdot(\nabla\cdot\Delta\sigma_{N})+\Delta\sigma_{N}:D(\Delta u_{N})\bigg]\ {\rm d}x=0.$ (49)
From (3.2), by using the $L^{2}$-estimate (32), the $H^{1}$-estimate (44), and the logarithmic Grönwall inequality in Lemma 2.2, we find the following bound:
$\sup_{t\in[0,T]}\biggl[\lVert u_{N}(t)\rVert_{H^{2}}^{2}+\lVert\sigma_{N}(t)\rVert_{H^{3}}^{2}\biggr]\leq C\exp(\exp(\exp(CT))),$ (50)
@@
Estimate on the time derivative and summary. Thanks to (50) and system (25), one can obtain directly that
$\sup_{t\in[0,T]}\biggl[\|\partial_{t}u_{N}(t)\|_{H^{2}}^{2}+\|\partial_{t}\sigma_{N}-\alpha^{2}\partial_{t}\Delta\sigma_{N}\|_{H^{1}}^{2}\biggr]\leq C\exp(\exp(\exp(CT))).$ (51)
This estimate follows because all the other terms (i.e. except the time derivative) in Eq. (25a) can be shown to lie in $C([0,T];H^{2}(\mathbb{T}^{2}))$, while all the other terms in Eq. (25d) are elements of $C([0,T];H^{1}(\mathbb{T}^{2}))$ (because of estimate (50)). Meanwhile, by directly applying integration by parts we deduce that

### 20 Mar 2026 v3

@@ 3.1 The Galerkin approximation scheme
$\partial_{t}u_{N}=\nabla\cdot\sigma_{N}+\mathbf{P}_{N}\mathcal{T}_{a}+\mathbf{P}_{N}\mathcal{T}_{w}^{N}+\Omega u^{\perp}_{N}-g\mathbf{P}_{N}\nabla H_{0},$ (3.1a)
$\begin{gathered}\frac{1}{\mathcal{E}}\partial_{t}\big(\sigma_{N}-\alpha^{2}\Delta\sigma_{N}\big)+\mathbf{Q}_{N}\bigg[\frac{4\mathcal{D}_{\epsilon}^{N}}{P}(\sigma_{N}-\frac{1}{2}\operatorname{Tr}\sigma_{N}\mathbb{I}_{2})\bigg]+\mathbf{Q}_{N}\bigg[\frac{\mathcal{D}_{\epsilon}^{N}}{2P}\operatorname{Tr}\sigma_{N}\mathbb{I}_{2}\bigg]\\ +\frac{\mathbf{Q}_{N}\mathcal{D}_{\epsilon}^{N}}{2}\mathbb{I}_{2}=D(u_{N}),\end{gathered}$ (3.1d)
$u\lvert_{t=0}=\mathbf{P}_{N}u_{0},\quad\sigma\lvert_{t=0}=\mathbf{Q}_{N}\sigma_{0},$ (3.1e)
where $\mathbf{P}_{N}$ is the projection of two-dimensional vector fields onto the Fourier modes up to order $N$, including the zeroth mode, while $\mathbf{Q}_{N}$ is the projection of $2\times 2$ (symmetric) matrices onto the Fourier modes up to order $N$. In system (3.1), we have used the notation
$\mathcal{T}_{w}^{N}$ $=c_{w}\rho_{w}\lvert U_{w}-u_{N}\rvert\bigg[(U_{w}-u_{N})\cos\theta+(U_{w}-u_{N})^{\perp}\sin\theta\bigg]$ (3.2)
@@
$\mathcal{D}_{\epsilon}^{N}$ $=\sqrt{\lvert D(u_{N})\rvert^{2}+\epsilon^{2}}.$ (3.3)
Notice that system (3.1) is a first-order system of ordinary differential equations for $\{\hat{u}_{k},\hat{\sigma}_{k}\}_{|k|\leq N}$, with a locally Lipschitz continuous vector field with respect to the unknowns. Thus, the Picard theory (see for example [28]) implies that, for each initial datum $(\mathbf{P}_{N}u_{0},\mathbf{Q}_{N}\sigma_{0})$, there exists a unique local-in-time solution.
@@ 3.2 Uniform-in-$(N,\epsilon)$ estimates
To establish the uniform-in-$N$ estimates for the solutions of system (3.1), let
$\tau_{N}\coloneqq\sigma_{N}+\frac{P}{2}\mathbb{I}_{2}.$ (3.4)
Recall that the pressure $P>0$ is constant. We then rewrite equation (3.1d) as follows:
$\begin{gathered}\frac{1}{\mathcal{E}}\partial_{t}\bigg(\tau_{N}-\alpha^{2}\Delta\tau_{N}\bigg)+\mathbf{Q}_{N}\bigg[\frac{4\mathcal{D}_{\epsilon}^{N}}{P}\bigg(\tau_{N}-\frac{1}{2}\operatorname{Tr}\tau_{N}\mathbb{I}_{2}\bigg)\bigg]\\ +\mathbf{Q}_{N}\bigg[\frac{\mathcal{D}_{\epsilon}^{N}}{2P}\operatorname{Tr}\tau_{N}\mathbb{I}_{2}\bigg]=D(u_{N}).\end{gathered}$ (3.5)
$L^{2}$-estimate. We start by computing the $L^{2}(\mathbb{T}^{2})$ estimates for the solution of system (3.1) on its maximal (positive) time interval of existence. Taking the $L^{2}$-inner product of (3.1a) and (3.5) with $u_{N}$ and $\tau_{N}$, respectively, applying integration by parts, and summing the results yields
$\begin{gathered}\frac{1}{2}\frac{\rm d}{\ {\rm d}t}\bigg[\lVert u_{N}\rVert_{L^{2}}^{2}+\mathcal{E}^{-1}\|\tau_{N}\|_{L^{2}}^{2}+\alpha^{2}\mathcal{E}^{-1}\|\nabla\tau_{N}\|_{L^{2}}^{2}\biggr]\\ +\int_{\mathbb{T}^{2}}\biggl[\frac{4\mathcal{D}_{\epsilon}^{N}}{P}|\tau_{N}-\frac{1}{2}\operatorname{Tr}\tau_{N}\mathbb{I}_{2}|^{2}+\frac{\mathcal{D}_{\epsilon}^{N}}{2P}|\operatorname{Tr}\tau_{N}|^{2}\biggr]\,\ {\rm d}{x}\\ =\int_{\mathbb{T}^{2}}\bigg[\mathcal{T}_{a}+\mathcal{T}_{w}^{N}+\Omega u^{\perp}_{N}-g\nabla H_{0}\bigg]\cdot u_{N}\ {\rm d}{x}\leq C+C\|u_{N}\|_{L^{2}}^{2},\end{gathered}$ (3.6)
@@
where we use the fact that $\cos\theta\geq 0$ since $\theta\in[0,\frac{\pi}{4}]$. Notice that the constant $C\in(0,\infty)$ in (3.6) depends only on $\Omega,g,P,H_{0},U_{a},U_{w},c_{a},c_{w},\phi,\theta,\rho_{a},$ and $\rho_{w}$, but not on $\epsilon$ and $N$. We remark that estimate (3.6) would also hold if one considers the intermediate model with the regularised and nonsimplified strain rate $\overline{\mathcal{D}}_{\epsilon}^{N}$ instead of $\mathcal{D}_{\epsilon}^{N}$ (as both have definite sign). Moreover, we also have used the following cancellation, which is similar in structure to the Oldroyd-B model,
$\int_{\mathbb{T}^{2}}\bigg[u_{N}\cdot(\nabla\cdot\sigma_{N})+\tau_{N}:D(u_{N})\bigg]\ {\rm d}{x}=0.$ (3.7)
Therefore, applying the Grönwall inequality to estimate (3.6) on a time interval of existence $[0,T]$ of the approximating Galerkin solution leads to
$\sup_{t\in[0,T]}\bigg[\lVert u_{N}(t)\rVert_{L^{2}}^{2}+\mathcal{E}^{-1}\left\lVert\sigma_{N}(t)+\frac{P}{2}\mathbb{I}_{2}\right\rVert_{L^{2}}^{2}+\alpha^{2}\mathcal{E}^{-1}\left\lVert\nabla\bigg[\sigma_{N}(t)+\frac{P}{2}\mathbb{I}_{2}\bigg]\right\rVert_{L^{2}}^{2}\bigg]\leq Ce^{CT},$ (3.8)
uniformly in $N$ (and also $\epsilon$). Since the right-hand side remains finite for any $T\in(0,\infty)$, we conclude that the time interval of existence of the approximating Galerkin solutions is $0,\infty)$ and estimate (3.8) therefore holds for any $T\in(0,\infty)$.
@@
$V_{w}:=U_{w}-u_{N}.$ (3.9)
Recalling (3.2), one can write
$\begin{gathered}-\int_{\mathbb{T}^{2}}\mathbf{P}_{N}\mathcal{T}_{w}^{N}\cdot\Delta u_{N}\ {\rm d}{x}=\sum_{i=1}^{2}\int_{\mathbb{T}^{2}}\mathbf{P}_{N}\partial_{i}\mathcal{T}_{w}^{N}\cdot\partial_{i}u_{N}\ {\rm d}{x}=\sum_{i=1}^{2}\int_{\mathbb{T}^{2}}\partial_{i}\mathcal{T}_{w}^{N}\cdot\partial_{i}u_{N}\ {\rm d}{x}\\ =\sum_{i=1}^{2}\int_{\mathbb{T}^{2}}\partial_{i}\mathcal{T}_{w}^{N}\cdot\partial_{i}{U_{w}}\ {\rm d}{x}-\sum_{i=1}^{2}\int_{\mathbb{T}^{2}}c_{w}\rho_{w}\bigl[\partial_{i}(\lvert V_{w}\rvert V_{w})\cos\theta\bigr]\cdot\partial_{i}V_{w}\ {\rm d}{x}\\ -\sum_{i=1}^{2}\int_{\mathbb{T}^{2}}c_{w}\rho_{w}\partial_{i}(|V_{w}|V_{w}^{\perp})\cdot\partial_{i}V_{w}\sin\theta\ {\rm d}{x}=:I_{1}+I_{2}+I_{3},\end{gathered}$ (3.10)
@@
$\partial_{i}(|V_{w}|V_{w})=|V_{w}|\partial_{i}V_{w}+\dfrac{V_{w}\cdot\partial_{i}V_{w}}{|V_{w}|}V_{w}.$ (3.11)
@@
$|I_{1}|\lesssim\|V_{w}\|_{L^{2}}\|\nabla V_{w}\|_{L^{2}}\lesssim 1+\|{u_{N}}\|_{L^{2}}^{2}+\|{\nabla u_{N}}\|_{L^{2}}^{2},$ (3.12)
@@
$=\left\lvert\sum_{i=1}^{2}\int_{\mathbb{T}^{2}}c_{w}\rho_{w}\dfrac{V_{w}\cdot\partial_{i}V_{w}}{|V_{w}|}V_{w}^{\perp}\cdot\partial_{i}V_{w}\sin\theta\ {\rm d}{x}\right\rvert$ (3.13)
$\leq c_{w}\rho_{w}|\sin\theta|\int_{\mathbb{T}^{2}}|V_{w}||\nabla V_{w}|^{2}\ {\rm d}{x}.$ (3.14)
Meanwhile, substituting (3.11) in $I_{2}$, one can compute
$I_{2}=-c_{w}\rho_{w}\cos\theta\int_{\mathbb{T}^{2}}\left(|V_{w}||\nabla V_{w}|^{2}+\frac{1}{4}\frac{\left|\nabla\big(|V_{w}|^{2}\big)\right|^{2}}{|V_{w}|}\right)\ {\rm d}{x}.$ (3.15)
Consequently, (3.10)–(3.15) leads to
$\int_{\mathbb{T}^{2}}\mathbf{P}_{N}\mathcal{T}_{w}^{N}\cdot\Delta u_{N}\ {\rm d}{x}=\int_{\mathbb{T}^{2}}\nabla\mathcal{T}_{w}^{N}:\nabla u_{N}\lesssim 1+\|{u_{N}}\|_{L^{2}}^{2}+\|{\nabla u_{N}}\|_{L^{2}}^{2}$ (3.16)
@@
Taking the $L^{2}$-inner product of (3.1a) and (3.1d) with $\Delta u_{N}$ and $\Delta\sigma_{N}$, respectively, applying integration by parts, and summing up the results leads to the following estimate
@@
$+\lVert\Delta\sigma_{N}\rVert_{L^{2}}^{2}+\lVert\nabla u_{N}\rVert_{L^{2}}^{2}+\|u_{N}\|_{L^{2}}^{2}+1$ (3.18)
where we have used (3.16), the Brézis-Gallouët-Wainger inequality (2.2), as well as Young’s inequality. Note that we have introduced a constant $R_{1}>0$ in order to make the quantity $R_{1}\sigma_{N}$ dimensionless (so that we can apply the Brézis-Gallouët-Wainger inequality), and in particular $R_{1}\sigma_{N}$ does not have physical units. One can choose $R_{1}=\lVert\sigma_{0}\rVert_{L^{\infty}}^{-1}$ for example (for nonzero initial data). We observe that for the ${H}^{1}$-estimate one can also treat the nonsimplified strain rate $\overline{\mathcal{D}}_{\epsilon}^{N}$ instead of $\mathcal{D}_{\epsilon}^{N}$ by using estimate (5.1). In addition, we have used the following cancellation identity
$\sum_{i=1}^{2}\int_{\mathbb{T}^{2}}\bigg[\partial_{i}u_{N}\cdot(\nabla\cdot\partial_{i}\sigma_{N})+\partial_{i}\sigma_{N}:D(\partial_{i}u_{N})\bigg]\ {\rm d}x=0.$ (3.19)
Thanks to the $L^{2}$-estimate in (3.8), we can apply the logarithmic Grönwall inequality from Lemma 2.2 to (3.18), where we set
@@
$\sup_{t\in[0,T]}\biggl[\|\nabla u_{N}\|_{L^{2}}^{2}+\mathcal{E}^{-1}\|\nabla\sigma_{N}\|_{L^{2}}^{2}+\alpha^{2}\mathcal{E}^{-1}\|\Delta\sigma_{N}\|_{L^{2}}^{2}\biggr]\leq C\exp(\exp(CT)),$ (3.20)
for some $C$ independent of $N,\epsilon$, and $T$, but depending on $\mathcal{E},\alpha$, and the initial data. Note that the double-exponential nature of the bound comes from the application of the logarithmic Grönwall inequality to the single-exponential $L^{2}$-estimate (3.8).
@@
$\nabla\mathcal{D}_{\epsilon}^{N}$ $=\frac{D(u_{N}):\big[\nabla D(u_{N})\big]}{\mathcal{D}_{\epsilon}^{N}}=\frac{\frac{1}{2}\nabla\lvert D(u_{N})\rvert^{2}}{\mathcal{D}_{\epsilon}^{N}},$ (3.21)
@@
In addition, we recall the following computation from [57,  p. 22-23], for $i,j=1,2$, with $V_{w}$ given in (3.9),
$\partial^{2}_{ij}\big(\lvert V_{w}\rvert V_{w}\big)=$ $\;\lvert V_{w}\rvert\partial^{2}_{ij}V_{w}+\frac{V_{w}\cdot\partial_{i}V_{w}}{\lvert V_{w}\rvert}\partial_{j}V_{w}+\frac{V_{w}\cdot\partial_{j}V_{w}}{\lvert V_{w}\rvert}\partial_{i}V_{w}$ (3.22)
@@
$\lesssim\big(\lVert V_{w}\rVert_{L^{\infty}}+\lVert V_{w}\rVert_{H^{1}}\big)\lVert V_{w}\rVert_{H^{2}},$ (3.23)
where we have applied the Ladyzhenskaya inequality (2.3).
Then, after taking the $L^{2}$-inner product of (3.1a) and (3.1d) with $\Delta^{2}u_{N}$ and $\Delta^{2}\sigma_{N}$, respectively, using integration by parts, and summing up the results, one obtains
@@
$+\|\sigma_{N}\|_{H^{2}}+1\bigg]\cdot\big[\lVert u_{N}\rVert_{H^{2}}^{2}+\lVert\sigma_{N}\rVert_{H^{3}}^{2}\big]+1,$ (3.24)
where we have used (3.23), the Brézis-Gallouët-Wainger inequality (2.2), the Sobolev embedding theorem as well as Young’s inequality. The terms on the second line are negative definite, and in the seventh line we estimate the term $\lVert\nabla^{2}u_{N}\rVert_{L^{2}}\lVert\nabla\sigma_{N}\rVert_{L^{4}}\lVert\Delta\sigma_{N}\rVert_{L^{4}}$ by means of the Sobolev embedding theorem, which gives $\lVert u_{N}\rVert_{H^{2}}\lVert\sigma_{N}\rVert_{H^{2}}\lVert\sigma_{N}\rVert_{H^{3}}$. Applying Young’s inequality leads to some of the terms on the last line of estimate (3.2). Due to the monotonicity of the logarithm, we can replace the arguments $\lVert\sigma_{N}\rVert_{H^{2}}$ and $\lVert u_{N}\rVert_{H^{2}}$ by their sum, which facilitates the application of the logarithmic Grönwall inequality.
@@
$\int_{\mathbb{T}^{2}}\bigg[\Delta u_{N}\cdot(\nabla\cdot\Delta\sigma_{N})+\Delta\sigma_{N}:D(\Delta u_{N})\bigg]\ {\rm d}x=0.$ (3.25)
From (3.2), by using the $L^{2}$-estimate (3.8), the $H^{1}$-estimate (3.20), and the logarithmic Grönwall inequality in Lemma 2.2, we find the following bound:
$\sup_{t\in[0,T]}\biggl[\lVert u_{N}(t)\rVert_{H^{2}}^{2}+\lVert\sigma_{N}(t)\rVert_{H^{3}}^{2}\biggr]\leq C\exp(\exp(\exp(CT))),$ (3.26)
@@
Estimate on the time derivative and summary. Thanks to (3.26) and system (3.1), one can obtain directly that
$\sup_{t\in[0,T]}\biggl[\|\partial_{t}u_{N}(t)\|_{H^{2}}^{2}+\|\partial_{t}\sigma_{N}-\alpha^{2}\partial_{t}\Delta\sigma_{N}\|_{H^{1}}^{2}\biggr]\leq C\exp(\exp(\exp(CT))).$ (3.27)
This estimate follows because all the other terms (i.e. except the time derivative) in equation (3.1a) can be shown to lie in $C([0,T];H^{2}(\mathbb{T}^{2}))$, while all the other terms in equation (3.1d) are elements of $C([0,T];H^{1}(\mathbb{T}^{2}))$ (because of estimate (3.26)). Meanwhile, by directly applying integration by parts we deduce that